# CLUE-Mark: Watermarking Diffusion Models using CLWE

### v3

@@ 3.1 Background on CLWE
Consider the scatter plot shown in Figure 3. If you were to look at the distribution of the points along either axis, it would look like a standard gaussian distribution. But if you were to plot the points along the line $y=-x$, you would notice that the points are periodic. That is, that they’re bunched together around $-1,-1/2,0,1/2,1,$ and so on. In two dimensions you can see this immediately by looking at the scatter plot, but what if the points are in three dimensions? The points would form a sphere from nearly any angle you look at, except if you found the “secret direction” you would notice the the points form slices and are again periodic [^1] The original CLWE presentation at https://youtu.be/9CB4KcoRB9U?t=60 has an animation that illustrates this effect clearly..
@@
This is the essence of the homogenous Continuous Learning With Errors (hCLWE) problem [^2] Homogenous CLWE is a special case of the more general CLWE problem.. Intuitively, it becomes harder to find the secret direction as you increase the number of dimensions, as the number of directions you have to check grows exponentially. Deciding whether a set of points (from now on let us call them samples) are normally distributed in all directions, or if there is a secret direction in which they are periodic has been proven to be as hard as solving worst-case lattice problems [4].
@@
To write this more formally, consider a unit vector $\mathbf{w}\in\mathbb{R}^{n}$ chosen uniformly at random, and a set of samples $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\in\mathbb{R}^{n}$ that are drawn from a standard gaussian distribution in all directions except for the direction $\mathbf{w}$, in which case the distribution is shown in Figure 4. This produces the well-known “gaussian-pancakes” distribution, which produces periodic slices with separation $\approx 1/\gamma$ and width $\approx\beta/\gamma$ in the secret direction $\mathbf{w}$, and standard gaussian in all others. The hCLWE problem is to determine the secret direction $\mathbf{w}$ given a polynomial number of samples $\{\mathbf{y}_{i}\}_{i\in[m]}$. The decision variant of hCLWE is to distinguish between samples from the hCLWE distribution and from the standard multivariate normal. This leads to Theorem 1, an informal statement of the CLWE hardness theorem; see [4] for the more formal statement, proof, and further details.
@@
Such lattice problems are considered cryptographically hard for appropriately chosen parameters $\gamma$ and $\beta$. Assuming that no PPT algorithm can solve them, then we can state that no PPT algorithm can solve the decision hCLWE problem non-negligible probability.
Figure 3: Example scatter plot of $10,000$ samples from the hCLWE distribution with secret direction $(1/\sqrt{2},-1/\sqrt{2})$, $\gamma=2,\beta=0.1$.
Figure 4: The hCLWE PDF with $\gamma=2,\beta=0.1$ in the secret direction compared with the standard gaussian distribution.
3.2 Watermarks from CLWE
Figure 5: CLUE-Mark watermark injection process.
@@ 3.2 Watermarks from CLWE
The hCLWE distribution assumes a number of high-dimensional samples, while the noise vector required for the initial diffusion model latents is a single vector. The solution is to divide the latent vector into a number of non-overlapping blocks, each of which is treated as a sample from the hCLWE distribution. The blocks must each be large enough to ensure undetectability, but small enough to ensure enough samples to recover the watermark even if some blocks are lost or corrupted by noise. As we will see in Section 6.1, for a latent vector of dimensions $4\times 64\times 64$ we choose a block size of either $4\times 1\times 1$ or $4\times 4\times 4$ for our watermark evaluations, but for the moment we leave this as an implementation parameter.
@@
In order to improve robustness to perturbations, we use a standard technique in image watermarking (see [15]): we apply the hCLWE samples in the frequency domain of the latents. That is we take the inverse discrete wavelet transform (IDWT) of the hCLWE samples to arrive at the latent vector used to generate the image. The process for image generation is shown in Figure 5. The watermark recovery process reverses this process, but requires additional steps in order to extract the watermark in the presence of noise and perturbations.
@@
We now have an approach that allows us to embed an hCLWE distribution in the latent space of a diffusion model, which we then use the produce an image. The question becomes: how do we determine if an image contains a watermark? From recent work [16, 17], we can invert the diffusion model to get an estimate of the original latent vector (assuming the image is not perturbed). Dividing the latent vector into blocks is then trivial, but the resulting blocks are subject to noise and perturbation. Given these noisy blocks and the secret direction of the hCLWE distribution, how do we determine if the watermark is present?
Figure 6: Watermark recovery process.
@@
$z_{i}=\gamma\langle\mathbf{y}_{i},\mathbf{w}\rangle\mod 1$
In the no-watermark case, we expect that for well-chosen parameters the $z_{i}$ values will be uniformly distributed in $[0,1)$. In the watermark case, we expect that the $z_{i}$ values will be concentrated near zero, with the strength of the concentration depending on $\beta$, noise, and perturbations. This produces the situation shown in Figure 7, which shows the result of a statistical simulation comparing $z_{i}$ values calculated from samples drawn from a standard normal, hCLWE, and hCLWE with noise from a normal distribution of width $0.2$.
Figure 7: Histogram of $|z_{i}|$ values from statistical simulation of $10,000$ samples.
Thus our problem boils down to distinguishing two distributions. There are many well-known distinguishing techniques one can use; we use the standard Kolmogorov-Smirnov (KS) statistic in this work. Intuitively, the KS statistic gives a measure of the difference between a set of samples from a reference distribution, with a higher score indicating that the samples are less likely to have been drawn that distribution. In our case we use the uniform distribution as the null hypothesis, and use KS to measure the distance of the given $z_{i}$ values from the uniform distribution. If the KS statistic is above a certain threshold, then we can reject the null hypothesis and say that the samples are not gaussian in the secret direction, and therefore the watermark is present.
In concrete terms, the KS statistic is the maximum distance between the empirical CDF of the samples and the expected CDF of the reference distribution. As illustrated in Figure 8, in the statistical simulation, the KS statistic for the standard normal samples is $0.036$, indicating a very small deviation from uniform, while the KS statistic for the hCLWE samples with $\beta=0.1$ is $0.774$ indicating a large deviation from uniform. Finally, the hCLWE samples with gaussian noise of variance $0.2$ produces a KS statistic of $0.575$, indicating a strong presence of the watermark despite the noise. Thus, using KS we get a strong indication of the presence of the watermark despite significant noise.
The resulting recovery process is shown in Figure 6. In terms of parameters, we set $n$ (the number of dimensions per CLWE sample), $\gamma$ and $\beta$ (CLWE parameters) for security of the reductions in Section 6.1, and $m$ follows from the size of the latent space.
Figure 8: Empirical CDF of CLWE errors from a statistical simulation of $10,000$ points drawn from several distributions.
4 Security Proof
@@ 4.1 Formal Definitions
$\rho_{s}(\mathbf{x})=\exp\left(-\pi\|\mathbf{x}/s\|^{2}\right)$ (1)
@@
No PPT distinguisher $D$ can solve the decision hCLWE problem with non-negligible advantage.
4.2 Undetectability
@@
Assuming the hCLWE assumption holds for $n$ dimensions with parameters $\gamma,\beta$, the CLUE-Mark watermarking scheme described in Section 3 with latent vectors of dimension $mn$ such that $m$ is at most polynomial in $n$ is undetectable by Definition 2.6.
@@
Clearly, if $D$ has advantage $\epsilon$ in distinguishing watermarked and unwatermarked images, then $D^{\prime}$ has at least $\epsilon$ advantage in distinguishing hCLWE samples from standard normal samples. Since all of the operations in $D^{\prime}$ are polynomial time, this contradicts the hCLWE assumption.
4.3 Sampling from hCLWE
Sampling from the hCLWE distribution is non-trivial. In [4] it is suggested that the hCLWE distribution be sampled by rejection-sampling from the standard normal distribution. However, this is not practical for our setup, as not only does this require drawing many more samples than needed, it also assumes the base distribution is the standard normal and easy to sample. In our case, the base distribution is the model’s latent space after a DWT transform. Rather than rejection sampling in this space, we propose a much simpler approach: beginning with samples of the appropriate base distribution (e.g. standard normal and then apply DWT), project the samples to the nearest CLWE point and then add noise in the secret direction. We can then reverse any transforms needed to get the samples in the required domain.
@@ 4       $z_{i}\leftarrow\mathcal{N}(0,\beta)$;
$\rho(\mathbf{y}^{\perp})\cdot\sum_{z\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta}(z)$
@@
$\rho(\mathbf{y}^{\perp})\cdot\sum_{z\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta/\gamma^{\prime}}\left(\tilde{y}^{\prime}-k\gamma/{\gamma^{\prime}}^{2}\right)$

### v4

@@ 3.1. Background on CLWE
Consider the scatter plot shown in Figure 3. If you were to look at the distribution of the points along either axis, it would look like a standard gaussian distribution. But if you were to plot the points along the line y=xy=-x, you would notice that the points are periodic. That is, that they’re bunched together around 1,1/2,0,1/2,1-1,-1/2,0,1/2,1, and so on. In two dimensions you can see this immediately by looking at the scatter plot, but what if the points are in three dimensions? The points would form a sphere from nearly any angle you look at, except if you found the “secret direction” you would notice that the points form slices and are again periodic [^2] The original CLWE presentation at https://youtu.be/9CB4KcoRB9U?t=60 has an animation that illustrates this effect clearly..
@@
This is the essence of the homogenous Continuous Learning With Errors (hCLWE) problem [^3] Homogenous CLWE is a special case of the more general CLWE problem.. Intuitively, it becomes harder to find the secret direction as you increase the number of dimensions, as the number of directions you have to check grows exponentially. Deciding whether a set of points (from now on let us call them samples) are normally distributed in all directions, or if there is a secret direction in which they are periodic has been proven to be as hard as solving worst-case lattice problems (Bruna et al., 2021).
@@
More formally, consider a unit vector $\mathbf{w}\in\mathbb{R}^{n}$ chosen uniformly at random, and a set of samples $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\in\mathbb{R}^{n}$ that are drawn from a standard gaussian distribution in all directions except for the direction $\mathbf{w}$, in which case the distribution is shown in Figure 4. This produces the well-known “gaussian-pancakes” distribution, which produces periodic slices with separation $\approx 1/\gamma$ and width $\approx\beta/\gamma$ in the secret direction $\mathbf{w}$, and standard gaussian in all others. The hCLWE problem is to determine the secret direction $\mathbf{w}$ given a polynomial number of samples ${\left\{\mathbf{y}_{i}\right\}}_{i\in[m]}$. The decision variant of hCLWE is to distinguish between samples from the hCLWE distribution and from the standard multivariate normal. This leads to Theorem 1, an informal statement of the CLWE hardness theorem; see (Bruna et al., 2021) for the more formal statement, proof, and further details.
@@
Such lattice problems are considered cryptographically hard for appropriately chosen parameters $\gamma$ and $\beta$. Assuming that no PPT algorithm can solve them, then we can state that no PPT algorithm can solve the decision hCLWE problem with non-negligible probability. Assumption 2.1 given later in Section 4.1 is the more precise version we rely on. It constraints the number of samples available.
Figure 3. Example scatter plot of $10,000$ samples from the hCLWE distribution with secret direction $(1/\sqrt{2},-1/\sqrt{2})$, $\gamma=2,\beta=0.1$.
Figure 4. The hCLWE PDF with $\gamma=2,\beta=0.1$ in the secret direction rescaled for comparison with a standard gaussian.
3.2. Creating Watermarks from CLWE
Figure 5. CLUE-Mark signal injection process showing how the latent vector is divided into blocks and transformed using the inverse discrete wavelet transform.
There are several technical challenges we overcome in making CLWE work reliably for diffusion models, which are explained next.
@@ 3.2. Creating Watermarks from CLWE
In order to use samples from the hCLWE distribution as inputs to a diffusion model, we must first overcome a small disparity: The hCLWE distribution assumes a number of high-dimensional samples, while the noise vector required for the initial diffusion model latents is a single vector. The solution is to divide the latent vector into a number of non-overlapping blocks, each of which is treated as a sample from the hCLWE distribution. The blocks must each be large enough to ensure undetectability, but small enough to ensure enough samples to recover the signal even if some blocks are lost or corrupted by noise. As we will see in Section 5.1, for a latent vector of dimensions $4\times 64\times 64$ we choose a block size of $2\times 4\times 4$ for our evaluations, but for the moment we leave this as an implementation parameter.
@@
Using hCLWE samples directly does not give robust recovery of the secret embedded. In order to improve robustness to perturbations and noise during recovery, we use a standard technique in image watermarking (see (Al-Haj, 2007)): we apply the hCLWE samples in the frequency domain of the latents, i.e., we take the inverse discrete wavelet transform (IDWT) of the hCLWE samples to arrive at the latent vector used to generate the image. The process for image generation is shown in Figure 5. The signal recovery process reverses this process, but requires additional steps in order to extract the signal in the presence of noise and perturbations.
@@
We now have an approach that allows us to embed an hCLWE distribution in the latent space of a diffusion model, which we then use to produce an image. The question becomes: how do we determine if an image contains a signal? From recent work (Hong et al., 2024; Song et al., 2020), we can invert the diffusion model to get an estimate of the original latent vector (assuming the image is not perturbed)—even without the original prompt or embedding vector used to create the image, albeit with some noise. Dividing the latent vector into blocks is then trivial, but the resulting blocks are subject to noise and perturbation. Given these noisy blocks and the secret direction of the hCLWE distribution, how do we determine if the secret signal is present?
Figure 6. Signal recovery process showing how the latent vector is extracted from the image, divided into blocks, and analyzed using the secret direction.
@@
$z_{i}=\gamma\langle\mathbf{y}_{i},\mathbf{w}\rangle\bmod 1$
In the no-signal case, we expect that for well-chosen parameters the $z_{i}$ values will be uniformly distributed in $[0,1)$. In the signal case, we expect that the $z_{i}$ values will be concentrated near zero, with the strength of the concentration depending on $\beta$, noise, and perturbations. This produces the situation shown in Figure 7, which shows the result of a statistical simulation comparing $z_{i}$ values calculated from samples drawn from a standard normal, hCLWE, and hCLWE with noise from a normal distribution of width $0.2$.
Figure 7. Rose diagram of $z_{i}$ values from statistical simulation of $10,000$ samples showing the distribution of errors for different cases, where a $z$ score of $1.0$ is mapped to $360\degree$.
We now need to decide between two hypotheseses: the null hypothesis is that the samples are from a standard normal distribution, and thus the $z_{i}$ values are uniformly distributed in $[0,1)$, while the alternative hypothesis is that the samples are from the hCLWE distribution, and thus the $z_{i}$ values have some a concentration around some (unimodal) peak. Fortunately, the Rayleigh test (Ruxton, 2017) is a well-known test in circular statistics that provides the optimal ability to distinguish between these two cases, and provides a $p$-value that indicates the probability that the null hypothesis is true given the observed data. Thus, if the $p$-value is below a certain threshold, we can reject the null hypothesis and conclude that the samples are from the hCLWE distribution, and thus the image contains a signal. For example, in the statistical simulation shown in Figure 7, if we restrict to the first 256 samples (as we would expect in our use case), the Rayleigh test gives a $p$-value of $36.5\%$ for the standard normal samples, $1.2\times 10^{-94}$ for the hCLWE samples, and $5.2\times 10^{-69}$ for the hCLWE samples with noise.
The resulting recovery process is shown in Figure 6. In terms of parameters, we set $n$ (the number of dimensions per CLWE sample), $\gamma$ and $\beta$ (CLWE parameters) for security of the reductions in Section 5.1, and $m$ is the size/dimensions of the latent vector divided by $n$.
4. Security Proof
@@ 4.1. Formal Definitions
(1) $\rho_{s}(\mathbf{x})=\exp\left(-\pi\|\mathbf{x}/s\|^{2}\right)$
@@
No PPT distinguisher $D$ can solve the decision hCLWE problem with non-negligible advantage when given a set of $m$ samples, where $m$ is at most polynomial in $n$.
The above assumption is stricter than the standard CLWE hardness, as it considers an adversary with a limited number of samples.
4.2. Undetectability
@@
Assuming the hCLWE assumption holds for $n$ dimensions with parameters $\gamma,\beta$, the CLUE-Mark scheme described in Section 3 with latent vectors of dimension mnmn such that $m$ is at most polynomial in $n$ is undetectable by Definition 6.
@@
If $D$ has advantage $\epsilon$ in distinguishing watermarked and unwatermarked images, then $D^{\prime}$ has at least $\epsilon$ advantage in distinguishing hCLWE samples from standard normal samples. Since all of the operations in $D^{\prime}$ are polynomial time, this contradicts the hCLWE assumption. ∎
4.3. Sampling from hCLWE
Sampling from the hCLWE distribution is non-trivial. In (Bruna et al., 2021) it is suggested that the hCLWE distribution be sampled by rejection-sampling from the standard normal distribution. However, this is not practical for our setup, as not only does this require drawing many more samples than needed, it also assumes the base distribution is the standard normal and easy to sample. In our case, the base distribution is the model’s latent space after a DWT transform. Rather than rejection sampling in this space, we propose a much simpler approach: beginning with samples of the appropriate base distribution (e.g. standard normal and then apply DWT), project the samples to the nearest CLWE point and then add noise in the secret direction. We can then reverse any transforms needed to get the samples in the required domain.
@@ 4 $z_{i}\leftarrow\mathcal{N}(0,\beta)$;
$\rho(\mathbf{y}^{\perp})\cdot\sum_{k\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta}(z)$
@@
$\rho(\mathbf{y}^{\perp})\cdot\sum_{k\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta/\gamma^{\prime}}\left(\tilde{y}^{\prime}-k\gamma/{\gamma^{\prime}}^{2}\right)$